# Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

### v6

@@ Abstract
I present a coordinate-free, symbolic framework for determining whether a given set of polygonal faces can form a closed, genus-zero polyhedral surface and for predicting how such a surface could be decomposed into internal tetrahedra. The method uses only discrete incidence variables, such as the number of internal tetrahedra $T$, internal gluing triangles $N_{i}$, and internal triangulation segments $S_{i}$, and applies combinatorial feasibility checks before any geometric embedding is attempted.
For polyhedra in normal form, I record exact incidence identities linking $V,E,F$ to a flatness parameter $S:=\sum_{f}(\operatorname{deg}f-3)$, and I identify parity-sensitive effects in $E$, $F$, and $S$. The external identities and parity-sensitive bounds hold universally for genus-0 polyhedral graphs. For internal quantities, I prove exact relations $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (with $S_{i}$ taken to be the number of interior edges) and obtain restricted linear ranges for internally decomposed polyhedra with the minimal number of added internal edges. Consequently, I propose a symbolic workflow that yields rapid pre-checks for structural impossibility, reducing the need for costly geometric validation in computational geometry, graphics, and automated modeling.
@@ 1 Introduction and Background
The approach is related in spirit to Ziegler’s symbolic treatment of polytopes [10] and to combinatorial efforts in tetrahedral decomposition (e.g., [4]), but differs in its exclusive reliance on integer logic and symbolic topological consistency. A classical result by Steinitz [8] shows that a graph is the 1-skeleton of a convex 3-polyhedron if and only if it is planar and 3-connected, providing the foundational combinatorial criterion for convex polyhedral graphs. While many known polyhedral validation tools rely on coordinate-based embeddings or optimization algorithms, the framework presented here derives topological invariants and structural bounds without any reference to spatial realization.
Recent extensions of Steinitz’s theorem have been explored for ball polyhedra [1], and Gallier and Quaintance [5] provide a contemporary combinatorial-topological survey covering shellings, Euler–Poincaré relationships, and triangulations. From a computational perspective, Below, De Loera, and Richter-Gebert [3] proved that finding a triangulation with the minimum number of tetrahedra is NP-complete. In contrast, the present method uses empirical observations to guide symbolic enumeration and narrow down the space of combinatorial configurations that may support enclosure and decomposition.
New contributions: A classical genus-zero identity (recorded here with a short derivation) for polyhedra in normal form is
$E=3V-6-S,\quad F=2V-4-S,$
where the flatness parameter
$S=\sum_{f}(\operatorname{deg}f-3)$
measures the total deviation from pure triangulation.
Combining these with the constraint that all vertices have degree at least $3$ yields sharp bounds:
Even $V$:
$0\leq S\leq\frac{3V}{2}-6,\quad\frac{3V}{2}\leq E\leq 3V-6,\quad\frac{V}{2}+2\leq F\leq 2V-4.$
Odd $V$:
$0\leq S\leq\frac{3V}{2}-\tfrac{13}{2}$
$\frac{3V+1}{2}\leq E\leq 3V-6,\quad\frac{V}{2}+\tfrac{5}{2}\leq F\leq 2V-4.$
All integer $S$ values within these ranges are conjecturally realizable (Heuristic Proposition 4.2). We give explicit families realizing many values (e.g., stacked triangulations followed by controlled coplanarity merges), and exhaustively confirm small $V$ cases, but I do not claim a full classification here. – For even $V$, this includes purely cubic graphs or pentagonal–trivalent graphs (e.g., cube, dodecahedron, tetrahedron). – For odd $V$, $S_{\max}$ occurs when all but one vertex are trivalent and the remaining vertex is 4-valent.
In all cases, $S$ subtracts exactly from a “triangulation budget” of 3V63V-6 edges.
A parity-sensitive behavior is observed: while Euler’s characteristic $\chi=V-E+F$ is invariant under $V$-parity changes, the structures of $E$, $F$, and $S$ differ in predictable ways between even and odd vertex counts.
@@
Finally, empirical enumeration of symbolic decompositions reveals additional internal constraints on $(T,N_{i},S_{i})$, as well as distribution-dependent effects such as “lettucing,” in which evenly distributed flatness changes achievable internal configurations compared to concentrated flatness.
@@
$a,b,c,\ldots$ Face types: triangle (3 sides), quadrilateral (4), pentagon (5), etc.
$a^{\prime},b^{\prime},c^{\prime},\ldots$ Vertex types: valency 3 (3 incident edges), valency 4, valency 5, etc.
@@ 3.1 Logical separation of results
    Proved genus-zero identities: These are exact, rigorously derived equalities and bounds that hold for all genus-zero polyhedra in normal form. Chief among them is the flatness identity
    $E=3V-6-S,\quad F=2V-4-S,$
    where $S=\sum_{f}(\operatorname{deg}f-3)$ measures the total deviation from triangulation. Interpreting $S=\sum_{f}(\operatorname{deg}f-3)$ as a flatness deficit, the identities imply the sharp ranges
    $\begin{cases}\dfrac{3V}{2},&V\ \text{even},\\[4.0pt] \dfrac{3V+1}{2},&V\ \text{odd}\end{cases}\;\;\leq\;E\;\;\leq\;3V-6,$
    $\begin{cases}\dfrac{V}{2}+2,&V\ \text{even},\\[4.0pt] \dfrac{V}{2}+\dfrac{5}{2},&V\ \text{odd}\end{cases}\;\;\leq\;F\;\;\leq\;2V-4,$
    $0\;\;\leq\;S\;\;\leq\;\begin{cases}\dfrac{3V}{2}-6,&V\ \text{even},\\[4.0pt] \dfrac{3V}{2}-\tfrac{13}{2},&V\ \text{odd}.\end{cases}$
    (Here the lower bound on $E$ uses $\operatorname{deg}(v)\geq 3$.) For each admissible $S$, $E$ and $F$ decrease by one from their triangulated values.
@@ 3.3 Surface flatness estimation
To quantify the flatness of a polyhedron’s surface, I introduce a discrete measure $S$, defined as the total number of triangulation segments added to convert all non-triangular external faces into triangles. Flatness in this context refers to the extent to which vertices on a given face are coplanar. The fewer triangular faces a polyhedron has initially, the more triangulation is required, and the higher its total flatness measure $S$.
Each polygonal face is triangulated by drawing non-intersecting segments between the nearest connectable vertices. The cumulative count of these added segments on all external faces constitutes the flatness measure $S$. When all faces of a polyhedron are triangles, $S=0$ and $F\equiv F_{\max}$. This discrete flatness measure is now linked directly to the proved genus-zero identity above, serving both as a predictor of external structure and as a measure of internal decomposition complexity.
@@ 3.4 Internal mesh construction from surface triangulation
Unlike the external surface of a polyhedron, which is both observable and abstractable with consistency, the internal structure must be inferred. Consequently, the number of tetrahedra required to tetrahedralize a polyhedron, such as a cube, depends critically on the configuration of face cuts. For example, as noted by Timalsina and Knepley [9], a cube may be divided into 5, 6, or 12 tetrahedra depending on the chosen strategy: five tetrahedra in an ideal face-aligned configuration, six using the marching tetrahedra algorithm, and twelve when a Steiner point is introduced at the cube’s center.
@@ 3.6 Tetrahedral decomposition and trivalent vertex prioritization
This approach differs from geometric decompositions which rely on slicing through faces or bisecting edges to create tetrahedral meshes. In contrast, the method presented here preserves polygonal integrity: despite triangulation, all faces remain whole, with fixed vertex sets; no edge bisection or midpoint insertion is permitted. Internal tetrahedra are constructed exclusively through explicit gluing at existing vertices, maintaining a combinatorially pure structure.
This work restricts attention to genus–0, normal–form decompositions built by peeling one outer shell of tetrahedra at a time, a procedure I will refer to informally as a shell–aligned ladder tetrahedralization (SALT). Each layer is created by coning existing boundary triangles and, only if necessary, introducing at most one new interior segment between boundary vertices to complete gluing across that layer. No interior vertices (Steiner points) and no edge bisections are allowed, and internal adjacencies occur only along triangular faces.
Within this scheme, a “minimal interior edge” (MIE) rule is observed: a new interior segment is introduced only when strictly necessary to complete the current layer; otherwise, none is added.
This SALT+MIE restriction excludes dense interior diagonals between boundary vertices (e.g. pulling triangulations of cyclic polytopes), which can create $\Theta(V^{2})$ tetrahedra without adding vertices. Our class supports the “ladder/shell” behavior analyzed in this paper.
@@ 4 Vertex-Driven External Polyhedral Properties: Intermediate Flatness Realizability and Parity Effects
I establish extremal bounds on external combinatorial quantities ($F$, $E$, and $S$) directly as functions of the vertex count $V$. These bounds emerge from Euler’s characteristic combined with flatness-based adjustments, and serve as the foundation for linking external structure to the internal decomposition results in later sections. This section proves that every integer flatness value in the admissible range $0\leq S\leq S_{\max}$ is realized by some genus-zero polyhedron in normal form. The realizability result is stated formally in Proposition 4.2.
@@ 4.1 Vertex Basis and Initial Angle Unit Calculation
Before stating the general formulas, I note two standard planar-graph facts for the maximally triangulated case ($S=0$):
  1. 1.
    Each face is a triangle, so $3F_{\max}=2E_{\max}$. Combining this with Euler’s formula $V-E_{\max}+F_{\max}=2$ gives:
    $F_{\max}=2(V-2),\quad E_{\max}=3(V-2).$
  2. 2.
    A coplanarity merge of two adjacent triangles into a polygonal face removes exactly one face and one edge, leaving $V$ unchanged. If $S$ counts the number of such merges from the maximally triangulated case, then:
    $F=F_{\max}-S,\quad E=E_{\max}-S.$
@@
For any simple, 3-connected planar graph on a genus-0 surface, with no digonal faces and in normal form with $V$ vertices, $E$ edges, $F$ faces, and flatness measure $S$ (the number of coplanarity merges from the maximally triangulated surface), the following identities hold:
@@
For a genus-0 polyhedron in normal form, consider first the triangulated case ($S=0$). In this case every face has degree $3$, so the incidence relation is
$3F_{\max}=2E_{\max}.$
Combining this with Euler’s formula $V-E_{\max}+F_{\max}=2$ yields
$F_{\max}=2V-4,\qquad E_{\max}=3V-6.$
Now, increasing the flatness parameter $S$ by one corresponds to replacing a triangular face by a face of degree $4$; this removes exactly one edge and one face from the triangulation. More generally, each unit of $S$ reduces both $E$ and $F$ by exactly one compared to the triangulated case. Therefore
$F=F_{\max}-S=2V-4-S,$
$E=E_{\max}-S=3V-6-S,$
as claimed. ∎
@@
In the symbolic flatness framework, both equality cases correspond to $S=S_{\max}$ and $F=F_{\min}$ (see Corollary 4.2). ∎
@@
$E_{\min}=\left\lceil\frac{3V}{2}\right\rceil=\frac{3V+1}{2}.$
@@
For odd $V$, the degree multiset achieving $E_{\min}$ necessarily has one 44–valent vertex and all others trivalent (by the handshake lemma). Realization for all odd $V$ is a separate existence question; examples exist for many $V$ (e.g., small explicit polyhedral graphs), but a full classification is beyond our scope.
@@
For $V\geq 4$, let $E_{\min}=\lceil 3V/2\rceil$. Then with the formulas from Proposition 4.1:
$S_{\max}=3V-6-E_{\min}=\begin{cases}\frac{3V}{2}-6,&\text{$V$ even},\\[2.0pt] \frac{3V}{2}-\tfrac{13}{2},&\text{$V$ odd},\end{cases}\qquad F_{\min}=2-V+E_{\min}=\begin{cases}\frac{V}{2}+2,&\text{$V$ even},\\[2.0pt] \frac{V}{2}+\tfrac{5}{2},&\text{$V$ odd}.\end{cases}$
Moreover, as $S$ increases by $1$ (flattening step), both $E$ and $F$ decrease by $1$, and for $k=0,1,\dots,S_{\max}$ one has E=3V6k,F=2V4k.E=3V-6-k,\;F=2V-4-k.
@@
On parity and the “half” constants. The halves (e.g., $S_{\max}=\tfrac{3V}{2}-\tfrac{13}{2}$ for odd $V$) come from the integer lower bound $E_{\min}=\lceil 3V/2\rceil$. Writing
$S_{\max}=3V-6-E_{\min}$
gives $S_{\max}=\tfrac{3V}{2}-6$ when $V$ is even and $S_{\max}=\tfrac{3V}{2}-\tfrac{13}{2}$ when $V$ is odd; these evaluate to integers at the respective parities. The corresponding formulas for $F_{\min}=2-V+E_{\min}$ have the same source.
@@
Substitute $E_{\min}$ into the formulas in Proposition 4.1. ∎

### v7

@@ Abstract
I present a coordinate-free, symbolic framework for determining whether a given set of polygonal faces can form a closed, genus-zero polyhedral surface and for predicting admissible internal tetrahedral decompositions consistent with incidence constraints. The method uses only discrete combinatorial variables, such as the number of tetrahedra $T$, internal gluing triangles $N_{i}$, and internal triangulation segments $S_{i}$, and applies feasibility checks prior to any geometric embedding.
For polyhedra in normal form, I record exact incidence identities linking $V$, $E$, and $F$ to a flatness parameter $S:=\sum_{f}(\deg f-3)$, and identify parity-sensitive extremal behavior in $E$, $F$, and $S$ arising from minimal vertex-degree constraints. These external identities and parity-dependent bounds hold for genus-zero polyhedral graphs under standard simplicity and connectivity assumptions. For internal quantities, I prove exact relations $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$, and derive restricted linear ranges for tetrahedral decompositions in normal form with no interior vertices. Together, these results yield a symbolic workflow for rapid pre-screening of combinatorially impossible configurations, reducing reliance on costly geometric validation in computational geometry, graphics, and automated modeling. An open-source R implementation is available at https://github.com/MoustaphaItani/polyenclose/.
@@ 1 Introduction and Background
The approach is related in spirit to Ziegler’s symbolic treatment of polytopes [ziegler2000] and to combinatorial efforts in tetrahedral decomposition (e.g., [eppstein2004]), but differs in its exclusive reliance on integer logic and symbolic topological consistency. A classical result by Steinitz [steinitz1922] shows that a graph is the 1-skeleton of a convex 3-polyhedron if and only if it is planar and 3-connected, providing the foundational combinatorial criterion for convex polyhedral graphs. While many known polyhedral validation tools rely on coordinate-based embeddings or optimization algorithms, the framework presented here derives topological invariants and structural bounds without any reference to spatial realization.
Recent extensions of Steinitz’s theorem have been explored for ball polyhedra [almohammad2020ballpolyhedra], and Gallier and Quaintance [gallier2022aspects] provide a contemporary combinatorial-topological survey covering shellings, Euler–Poincaré relationships, and triangulations. From a computational perspective, Below, De Loera, and Richter-Gebert [below2004complexity] proved that finding a triangulation with the minimum number of tetrahedra is NP-complete. In contrast, the present method uses empirical observations to guide symbolic enumeration and narrow down the space of combinatorial configurations that may support enclosure and decomposition.
New contributions. For genus-zero polyhedral graphs in normal form, I record exact external identities linking $(V,E,F)$ to the face-degree excess (flatness)
$S:=\sum_{f}(\deg f-3),$
namely
$E=3V-6-S,\qquad F=2V-4-S,$
proved in Proposition 4.1. Combining these with the minimum degree condition $\deg(v)\geq 3$ yields parity-sensitive extremal bounds on $E$, $F$, and $S$ (Lemma 4.1 and Corollary 4.2).
The resulting extremal ranges are summarized in Corollary 4.2.
Remark.
For even $V$, equality in Lemma 4.1 is achieved by cubic (polyhedral) graphs such as prisms, the cube, and the dodecahedron. For odd $V$, equality requires exactly one 44–valent vertex and all others trivalent; explicit realizations exist for many $V$, though no classification is claimed here.
@@
Finally, empirical enumeration suggests a sharp rigidity phenomenon at the top of the internal ladder: bipyramidal (bipyramidal–prism) polyhedra appear to be the only normal–form surfaces attaining $T_{\max}$, and their internal invariants $(T,N_{i},S_{i})$ are essentially SS–inflexible, admitting no variation in the external parameters $(E,F,S)$ that is compatible with a valid tetrahedral decomposition. More generally, determining how the realized set of admissible triples $(T,N_{i},S_{i})$ contracts as the surface flatness $S$ increases remains open, and may admit either an empirical resolution via systematic enumeration or a theoretical one.
@@
$F_{3},F_{4},F_{5},\ldots$ Face types: triangle (3 sides), quadrilateral (4), pentagon (5), etc.
$V_{3},V_{4},V_{5},\ldots$ Vertex types: valency 3 (3 incident edges), valency 4, valency 5, etc.
@@ 3.1 Logical separation of results
    Proved genus-zero identities: These are exact, rigorously derived equalities and bounds that hold for all genus-zero polyhedra in normal form. Chief among the proved results is the flatness identity
    $E=3V-6-S,\qquad F=2V-4-S,$
    where $S:=\sum_{f}(\deg f-3)$ (Proposition 4.1). Combining this with the minimum degree condition $\deg(v)\geq 3$ yields parity-sensitive extremal bounds on $E$, $F$, and $S$ (Lemma 4.1 and Corollary 4.2).
@@ 3.3 Surface flatness estimation
To quantify flatness on the surface, I use the face-degree excess
$S:=\sum_{f}(\deg f-3),$
which equals the total number of non-intersecting diagonals needed to triangulate all polygonal faces. Thus $S=0$ if and only if every face is triangular, and increasing $S$ corresponds to fewer faces and edges relative to the triangulated case via Proposition 4.1.
Remark.
Because the embedding is planar with all faces homeomorphic to disks (i.e., a 22–cell embedding of $S^{2}$), each edge is incident to exactly two distinct faces.
Each polygonal face is triangulated by drawing non-intersecting segments between the nearest connectable vertices. The cumulative count of these added segments on all external faces constitutes the flatness measure $S$. When all faces of a polyhedron are triangles, $S=0$ and $F\equiv F_{\mathrm{ub}}$.
@@ 3.4 Internal mesh construction from surface triangulation
Unlike the external surface of a polyhedron, which is both observable and abstractable with consistency, the internal structure must be inferred. Consequently, the number of tetrahedra required to tetrahedralize a polyhedron, such as a cube, depends critically on the configuration of face cuts. For example, as noted by Timalsina and Knepley [timalsina2023tetrahedralization], a cube may be divided into 5, 6, or 12 tetrahedra depending on the chosen strategy: five tetrahedra in an ideal face-aligned configuration, six using the marching tetrahedra algorithm, and twelve when a Steiner point is introduced at the cube’s center.
@@ 3.6 Tetrahedral decomposition and trivalent vertex prioritization
This approach differs from geometric decompositions which rely on slicing through faces or bisecting edges to create tetrahedral meshes. In contrast, the method presented here preserves polygonal integrity: despite triangulation, all faces remain whole, with fixed vertex sets. No edge bisection is allowed and no midpoint or interior vertex insertion (Steiner points) is permitted. As internal tetrahedra are constructed exclusively through explicit gluing at existing vertices, internal adjacencies occur only along triangular faces.
This work restricts attention to genus–0, normal–form decompositions built by peeling one outer shell of tetrahedra at a time, a procedure I will refer to informally as a shell–aligned ladder tetrahedralization (SALT). Each layer is created by coning existing boundary triangles. Within this scheme, a “minimal interior edge” (MIE) rule is observed: a new interior segment between boundary vertices is introduced only when strictly necessary to complete the current triangular gluing layer; otherwise, none is added. This SALT+MIE restriction excludes dense interior diagonals between boundary vertices (e.g. pulling triangulations of cyclic polytopes), which can create $\Theta(V^{2})$ tetrahedra without adding vertices.
@@ 4 Vertex-Driven External Polyhedral Properties: Intermediate Flatness Realizability and Parity Effects
I establish extremal bounds on external combinatorial quantities ($F$, $E$, and $S$) directly as functions of the vertex count $V$. These bounds emerge from Euler’s characteristic combined with flatness-based adjustments, and serve as the foundation for linking external structure to the internal decomposition results in later sections. This section establishes the admissible parity-sensitive bounds on $E$, $F$, and $S$ and presents heuristic evidence that every integer flatness value $0\leq S\leq S_{\mathrm{ub}}$ is realizable by some genus-zero polyhedron in normal form. The realizability result is stated formally in Proposition 4.3.
@@ 4.1 Vertex Basis and Initial Angle Unit Calculation
Remark (Triangulated baseline and flattening intuition).
When the boundary is fully triangulated (i.e. $S=0$), one has $F_{\mathrm{ub}}=2V-4$ and $E_{\mathrm{ub}}=3V-6$. Interpreting $S=\sum_{f}(\deg f-3)$ as the total triangulation deficit, each unit increase of $S$ corresponds to removing one diagonal from a triangulated refinement, decreasing both $E$ and $F$ by one. The formal identities are stated and proved in Proposition 4.1.
@@
For a genus-0 polyhedral graph in normal form with $V$ vertices, $E$ edges, $F$ faces, and face-degree excess
$S:=\sum_{f}(\deg f-3),$
the following identities hold:
@@
Because the embedding is a 2–cell embedding of $S^{2}$, each edge is incident to exactly two faces, hence $\sum_{f}\deg f=2E$. By definition, $S=\sum_{f}(\deg f-3)=2E-3F$. Euler’s formula gives F=EV+2F=E-V+2. Substituting,
S=2E3(EV+2)=3V6E,S=2E-3(E-V+2)=3V-6-E,
so E=3V6SE=3V-6-S, and then F=EV+2=2V4SF=E-V+2=2V-4-S. ∎
@@
In the symbolic flatness framework, both equality cases correspond to $S=S_{\mathrm{ub}}$ and $F=F_{\mathrm{lb}}$ (see Corollary 4.2). ∎
@@
$E_{\mathrm{lb}}=\left\lceil\frac{3V}{2}\right\rceil=\frac{3V+1}{2}.$
@@
For odd $V$, the degree multiset achieving $E_{\mathrm{lb}}$ necessarily has one 44–valent vertex and all others trivalent (by the handshake lemma). Realization for all odd $V$ is a separate existence question; examples exist for many $V$ (e.g., small explicit polyhedral graphs), but a full classification is beyond our scope.
@@
For $V\geq 4$, let $E_{\mathrm{lb}}=\lceil 3V/2\rceil$. Then with the formulas from Proposition 4.1:
$S_{\mathrm{ub}}=3V-6-E_{\mathrm{lb}}=\begin{cases}\frac{3V}{2}-6,&\text{$V$ even},\\[2.0pt] \frac{3V}{2}-\tfrac{13}{2},&\text{$V$ odd},\end{cases}\qquad F_{\mathrm{lb}}=2-V+E_{\mathrm{lb}}=\begin{cases}\frac{V}{2}+2,&\text{$V$ even},\\[2.0pt] \frac{V}{2}+\tfrac{5}{2},&\text{$V$ odd}.\end{cases}$
Moreover, as $S$ increases by $1$ (flattening step), both $E$ and $F$ decrease by $1$, and for $k=0,1,\dots,S_{\mathrm{ub}}$ one has E=3V6k,F=2V4k.E=3V-6-k,\;F=2V-4-k.
@@
On parity and the “half” constants. The halves (e.g., $S_{\mathrm{ub}}=\tfrac{3V}{2}-\tfrac{13}{2}$ for odd $V$) come from the integer lower bound $E_{\mathrm{lb}}=\lceil 3V/2\rceil$. Writing
$S_{\mathrm{ub}}=3V-6-E_{\mathrm{lb}}$
gives $S_{\mathrm{ub}}=\tfrac{3V}{2}-6$ when $V$ is even and $S_{\mathrm{ub}}=\tfrac{3V}{2}-\tfrac{13}{2}$ when $V$ is odd; these evaluate to integers at the respective parities. The corresponding formulas for $F_{\mathrm{lb}}=2-V+E_{\mathrm{lb}}$ have the same source.
@@
Substitute $E_{\mathrm{lb}}$ into the formulas in Proposition 4.1. ∎
This proposition is included only to formalize the triangulated baseline and the “flattening ladder” interpretation; the definitive identities are Proposition 4.1 and Corollary 4.2.
Proposition 4.2 (Triangulated extrema and invariance of $\chi$).
For a genus-zero polyhedral surface composed entirely of triangles,
$F_{\mathrm{ub}}=2(V-2),\qquad E_{\mathrm{ub}}=3(V-2).$
$F(S)=F_{\mathrm{ub}}-S,\qquad E(S)=E_{\mathrm{ub}}-S,\qquad V-E(S)+F(S)=2.$
Proof.
Each triangle has three edges and each edge is shared by two triangles, so $2E_{\mathrm{ub}}=3F_{\mathrm{ub}}$. Substituting $E_{\mathrm{ub}}=\tfrac{3}{2}F_{\mathrm{ub}}$ into Euler’s formula $V-E_{\mathrm{ub}}+F_{\mathrm{ub}}=2$ yields $F_{\mathrm{ub}}=2(V-2)$ and $E_{\mathrm{ub}}=3(V-2)$. A coplanarity merge of two adjacent triangles removes one surface edge and one triangle, so $F\mapsto F-1$ and $E\mapsto E-1$ with $V$ unchanged. Hence VE+FV-E+F is invariant and equal to $2$ for all admissible $S$. ∎
Equivalent parameterization for fixed $V$ from the lower extremum.
Equivalently, indexing the flattening ladder from the lower extremal configuration $(S_{\mathrm{ub}},E_{\mathrm{lb}},F_{\mathrm{lb}})$, one may write
$S_{k}=S_{\mathrm{ub}}-k,\qquad E_{k}=E_{\mathrm{lb}}+k,\qquad F_{k}=F_{\mathrm{lb}}+k,$
for $k=0,1,\dots,S_{\mathrm{ub}}$. This parameterization enumerates all admissible external configurations consistent with the flattening rule and Euler invariance.